# Fast, Stable and Efficient Approximation of Multi-parameter Persistence Modules with MMA

### v3

@@ 5.2.2 Vineyards algorithm and matching
First, note that all maps in that diamond are either injective of corank 1 or surjective of nullity 1 (since they are all induced by adding a positive or negative chain of the corresponding reduced bases). Moreover, by the Mayer-Vietoris theorem, the following sequence is exact:
$N_{\text{id},i-1}=N_{(i\,i+1),i-1}$ $=H_{*}(C^{\text{id}}_{i-1})\rightarrow N_{(i\,i+1),i}\oplus N_{\text{id},i}=H_{*}(C^{(i\,i+1)}_{i})\oplus H_{*}(C^{\text{id}}_{i})$
$\xrightarrow{(x,y)\mapsto y-x}H_{*}(C^{\text{id}}_{i+1})=N_{\text{id},i+1}=N_{(i\,i+1),i+1}.$
Such diamonds are called transposition diamonds, and it has been shown in [MO15, Theorem 2.4] that the morphism induced by $\sigma_{\mathrm{vine}}^{i}$ between the lower and upper parts of the diamond matches bars with same representative positive chains together. Hence, bars in ${\mathcal{B}}(N_{\rm id})$ and ${\mathcal{B}}(N_{(i\,i+1)})$, and thus the bars in ${\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l_{1}}})$ and ${\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l_{2}}})$ are matched under $\sigma_{\mathrm{vine}}$ are associated to the same summand of ${\mathbb{M}}$, and thus are compatible. In particular, when ${\mathbb{M}}$ is decomposable into intervals summands, $\sigma_{\mathrm{vine}}$ is exact. Finally, by repeating this argumentation with the other transpositions induced by the simplices reordering from $l_{i}$ to $l_{i+1}$, for $i=2,\dots,k-1$ one has that the same is true for bars in ${\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l}})$ and ${\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l^{\prime}}})$. ∎
@@ 6 Numerical experiments
In this section, we showcase the performances of MMA. More precisely, we empirically study how the output quality and running time depend on the precision $\delta$ and the number of lines $|L|$. Then, we compare the running times of MMA with those of Rivet [LW15] and the elder-rule staircode (ERS) [CKMW20]. Finally, we investigate how running time is affected by the number of filtrations. All experiments were done on a laptop with AMD Ryzen 4800 CPU and 16GB of RAM. Our code is publicly available at https://github.com/DavidLapous/multipers.
@@
In our first two experiments, we focus on two real-world data sets of point clouds. The first ones, called LargeHypoxicRegion, were obtained from immunohistochemistry in [Vip20b]. These are made of a few thousand points, each representing a single cell. The others were obtained from applying time-delay embedding in $\mathbb{R}^{2}$ on time series taken from a few data sets (Wine, Plane, OliveOil, Coffee) from the UCR archive [Che15]. On all of these data sets, we computed bi-filtrations using the standard Vietoris-Rips filtration, and the superlevel sets of a Gaussian kernel density estimation (with bandwidth parameter $0.1d$ where $d$ is the diameter of the dataset), and we applied MMA and its competitors on the corresponding multi-parameter persistence modules in homology dimensions 00 and $1$ (note that the ERS can only compute dimension 00).
In our third experiment, we measure the dependence on the number of filtrations using a synthetic data set obtained by sampling $300$ points in the unit square $[0,1]^{2}$, computing their Alpha simplicial complex, and generating $n$ random filtration values on the points.
Finally, it is worth noting that we used multi-parameter edge collapses [AKP22] in order to simplify the multi-parameter persistence modules (without losing information) as much as possible before applying Rivet and MMA. The timing of this simplifications are not taken into account, but they are not the computational bottleneck of our computations. Furthermore, for the LargeHypoxicRegion, we thresholded the Rips edges at $0.02$, leading to simplicial complexes of $\sim 85k$ and $\sim 125k$ simplices respectively, after simplifications.
@@
We first empirically validate Propositions 3.4 and 4.13 by measuring how far is the output of MMA from the data when the precision parameter $\delta$ decreases and the number of lines in $L$ increases. For this, we used the bottleneck distances between the fibered barcodes (on $100$ random diagonal lines) of the outputs of MMA and the ones of the underlying modules as a proxy for the interleaving distances (since they are practically very difficult to evaluate). We call this the estimated matching distance. Results are displayed in Figure 13. One can see that the convergence is empirically linear with the number of lines $|L|$, which is in line with Propositions 3.4 and 4.13 (since $|L|$ increases linearly as $\delta$ decreases for a fixed $n$). Note that the distance even reaches 00 on a few cases.
@@
We now compare the running times of MMA with those of Rivet and the ERS. Results can be found in Table 1. It is worth noting that on several occasions, Rivet and the ERS could not produce outputs in reasonable time, due (among other things) to large memory consumption. On the other hands, the fact that MMA produces discretely presented intervals allows to encode them in a sparse manner with their corners. Note that computations with Rivet can be also approximated by coarsening the filtration values, and thus the module. In practice, this corresponds to restricting the 2-module to a $\kappa=\kappa_{x}\times\kappa_{y}$ grid, where $\kappa_{x},\kappa_{y}\in\mathbb{N}$ are the resolutions in both axes, see [LW15, Section 8.2]. The parameters $\kappa$ and $|L|$ have (roughly) the same role, and can be related with $\sqrt{\kappa}\simeq|L|$ (for a given, prescribed precision). Overall, we find that the running times of MMA outperforms its competitors, except when $\kappa$ is very small. However, in this case, the corresponding output of Rivet is very crude as the module is restricted on just a few points, whereas MMA produces intervals that are accurate along whole straight lines: we find that for large data sets, MMA is the only method that can produce accurate representations.
@@
Coffee $0.21\pm 0.01,0.18\pm 0.01$ $9.75\pm 5.92,0.35\pm 0.12$ ,0.95±0.56--,0.95$\pm$ 0.56- - , 0.95 ± 0.56 $0.34\pm 0.04$ $0.0093\pm 0.001$ $0.024\pm 0.001$ $0.16\pm 0.008$
Plane $0.19\pm 0.005,0.18\pm 0.03$ $4.36\pm 2.24,0.28\pm 0.04$ $33.3\pm 17.5,0.56\pm 0.17$ $0.09\pm 0.03$ $0.004\pm 0.0$ $0.012\pm 0.001$ $0.095\pm 0.004$
Wine $0.21\pm 0.003,0.19\pm 0.007$ $8.50\pm 2.00,0.22\pm 0.01$ ,0.28±0.023--,0.28$\pm$ 0.023- - , 0.28 ± 0.023 $0.22\pm 0.04$ $0.004\pm 0.0$ $0.016\pm 0.0$ $0.129\pm 0.002$
OliveOil $0.21\pm 0.004,0.19\pm 0.002$ $5.55\pm 1.20,0.31\pm 0.016$ ,0.82±0.17--,0.82$\pm$ 0.17- - , 0.82 ± 0.17 $1.39\pm 0.03$ $0.026\pm 0.001$ $0.058\pm 0.001$ $0.37\pm 0.006$
Worms $0.29\pm 0.082,0.23\pm 0.23$ $19.9\pm 14.4,4.60\pm 5.0$ ,31.36±36.24--,31.36$\pm$ 36.24- - , 31.36 ± 36.24 $3.85\pm 0.1$ $0.22\pm 0.11$ $0.34\pm 0.15$ $1.35\pm 0.41$
LargeHypoxicRegion1 $1.73,2.88$ ,234--,234- - , 234 ,--,--- - , - - --- - $26.4$ $26.6$ $59.4$
LargeHypoxicRegion2 $2.39,6.04$ ,--,--- - , - - ,--,--- - , - - --- - $57.3$ $54.3$ $102.9$
@@
Interestingly, computing 0-dimensional homology is sometimes slower than 1-dimensional homology for Rivet; as it relies on computing minimal free presentations, we think that this comes from the fact that minimal presentations in homology dimension 0 can be more complex than their counterparts in homology dimension 1, i.e., they have much more generators. We also investigate how running times of MMA depend on the number of lines. Unsurprisingly, one can see from Figure 14 that running time increases with the number of lines. However, the dependency looks empirically sublinear, which could come from the fact that even though there are more lines, these lines are closer to each other, and thus matching them with vineyards requires fewer computation steps. This is also highlighted by the running times of LargeHypoxicRegion2, Table 1 which are smaller when computing it over $1~{}000$ lines than $100$ lines.
[Uncaptioned image] Figure 14: Running time (s) needed to run MMA on the UCR datasets.
[Uncaptioned image] Figure 15: Running time (s) of MMA w.r.t. the number $n$ of filtrations.
@@
Finally, we investigate how the running times of MMA depend on the number $n$ of filtrations. While most of the approaches in the literature are limited to $n=2$ parameters, one can see from Figure 15 that MMA can produce outputs in a reasonable amount of time for up to $n\simeq 10$ parameter filtrations. As expected from the complexity of MMA in Equation 8, the running times increase exponentially with $n$.
@@ 7 Conclusion
In this article, we presented MMA: a new scheme for computing topological invariants (in the form of candidate decompositions) of multi-parameter persistence modules, in a fast and robust way, and whose complexity, running time, and approximation error can be controlled by user-defined parameters. We then showcased the performances of our method on synthetic and real data sets. Our code is publicly available at https://github.com/DavidLapous/multipers.
There are a number of interesting directions for future work.
@@
    We conjecture that the family of candidate decompositions produced by varying matching functions completely recovers multi-dimensional persistence modules.
@@
    We proved that our candidate has bounded approximation error when approximating interval decomposable modules. We also proved stability even when the underlying module is not interval decomposable. Is there some kind of optimality result describing the relationship of the candidate decomposition to the underlying module in this case?
@@
    Recently, stability results have become available for specific multi-parameter persistence modules [BL20]. We expect that these can be used to infer theoretical confidence regions and convergence rates for our candidate decompositions.

### v4

@@ 6.2.2 Vineyards algorithm and matching
Then, using the fact that any cyclic permutation $\sigma_{i_{1},\dots,i_{k}}=(i_{1}\,i_{2})\circ(i_{2}\,i_{3})\circ\cdots(i_{k-1}\,i_{k})$ is a product of transpositions, it follows that every permutation is a product of transpositions as well. Finally, if $(i,j)$ is a transposition with $1\leq i<j\leq n$, using the following relation:
$(i\,j)=(i\,i+1)\circ\cdots\circ(j-1\,j)\circ(j-2\,j-1)\circ\cdots\circ(i\,i+1),$
one can see that every permutation is in fact the product of adjacent transpositions. In practice, this product can be retrieved using a sorting algorithm, such as the insertion sort, or the bubble sort (see, e.g., (42)).
Remark 15 (Lazy vineyard update).
Note that the (ii+1)(i\,i+1) swaps can be directly inferred with the approach provided in (38, Section 4, Lazy minimazation). More formally, using the same notations, the swaps (ii+1)(i\,i+1) only occur precisely when two incomparable filtration values satisfy $x_{i}^{l}=x_{i+1}^{l}$ for some line $l$, which, in the two-parameter case, corresponds to the presence of a line crossing the point $(\max\left\{(x_{i})_{1},(x_{i+1})_{1}\right\},\max\left\{(x_{i})_{2},(x_{i+1})_{2}\right\})$. This suggests that our MMA algorithm can be trivially extended to fibered barcodes involving only such lines (instead of $\delta$-grids of lines), thus reducing its running time. We stick to grids in this article for the sake of clarity.
Proposition 17 (Vineyards is compatible for $n=2$).
Let $\mathbb{M}$ be a $2$-parameter persistence module, and $L$ be an ordered set of diagonal lines $L:=(l_{i})_{1\leq i\leq N}$ with increasing basepoints. [^6] Ordering $L$ in such a way is possible precisely because $n=2$. For each index $1\leq i\leq N$, let $\sigma_{i}$ be an arbitrary compatible matching function between ${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l_{i}}}\right)$ and ${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l_{i+1}}}\right)$ (obtained with, e.g., Proposition 16). Then, the matching function $\sigma=(\sigma_{i})_{1\leq i\leq N}$ is a compatible matching function on $L$.
Proof.
First, note that, given a compatible matching function $\sigma$ between two diagonal lines $l,l^{\prime}\in L$, and letting $b\in{\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right),b^{\prime}\in{\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l^{\prime}}}\right)$ be any non-trivial pair of matched bars, then, assuming that, e.g., $(x,0)\leq(x^{\prime},0)$ are the basepoints of $l$ and $l^{\prime}$ respectively, one has $\min(b)_{1}\leq\min(b^{\prime})_{1}$ and $\max(b)_{2}\geq\max(b^{\prime})_{2}$. See Figure 17.
Figure 17: Let $l_{1}<l_{2}$ be two diagonal lines of $\mathbb{R}^{2}$, and $x$, $y$ be two matched points in $l_{1}$ and $l_{2}$ respectively. Then, if the matching is compatible, one must have $x_{1}\leq y_{1}$ and $x_{2}\geq y_{2}$.
Now, let $x\in l_{1}$ be the grade of a given bar endpoint. A point $y$ is strictly incomparable with $x$ if either:
  1. 1.
    $y_{1}\leq x_{1}$ and $y_{2}\geq x_{2}$, or
  2. 2.
    $y_{1}\geq x_{1}$ and $y_{2}\leq x_{2}$.
Now, assuming that $y\in l_{2}$ and $x$ and $y$ are strictly incomparable, Case 1 can be excluded. Indeed, in that case, there exist constants $x_{0},y_{0},s,t\in\mathbb{R}$ such that $x=(x_{0}+s,s)$ and $y=(y_{0}+t,t)$, and hence:
[…]
This shows that two bar endpoints matched by a compatible matching function between two consecutive diagonal lines with increasing basepoints must satisfy Case 2, and thus, that a sequence of compatible matching functions is still compatible. ∎
@@ 7 Numerical experiments
In this section, we showcase the performances of MMA. More precisely, we empirically study how the output quality and running time depend on the precision $\delta$ and the number of lines $|L|$. Then, we compare the running times of MMA with those of Rivet (47) and the elder-rule staircode (ERS) (17), which are our closest competitors in terms of producing visual and interpretable descriptors of persistence modules. Finally, we investigate how running time is affected by the number of filtrations. All experiments were done on a laptop with AMD Ryzen 4800 CPU and 16GB of RAM. Our code is part of the multipers library (50) and is publicly available at https://github.com/DavidLapous/multipers.
Interpretation.
We first qualitatively show how to interpret our candidate decompositions on a toy dataset in Figure 18, using two different two-parameter filtrations (Čech and (edge-collapsed) Rips with sublevel sets of codensity) on a point cloud, computed using (2). In these examples, we consider a coordinate $x\in\mathbb{R}^{2}$ that is included in the support of some specific summands of the candidate decomposition, i.e., in some specific colored shapes in the plot. Then, we look back at the filtration at this coordinate $x$ and we identify the corresponding cycles. Note that, at each of these points $x$, the cycle representatives (from Theorem 15) are already calculated when running MMA and can thus be used for interpretation without additional computational cost.
((a)) (Left) The Čech complex, at radius $0.2$, of the points with codensity values larger than $3$. (Right) The corresponding candidate decomposition, with a red dot at the coordinates fixed by the radius and codensity values used on the left.
((b)) (Left) The (edge-collapsed) Rips complex, at radius $0.4$, of the points with codensity values larger than $3$. (Right) The corresponding candidate decomposition, with a red dot at the coordinates fixed by the radius and codensity values used on the left.
Figure 18: Interpretation of candidate decompositions computed by MMA on a toy dataset.
We also provide another example, containing $25,000$ points uniformly sampled on the unit square (noise), and $25,000$ points sampled on three distinct annuli with different sizes concentration levels (signal). See Figure 19. As before, we then compute the Čech-codensity filtration on this point cloud. The first cycle is very dense, so it should appear quickly w.r.t. the codensity filtration (i.e., the lower part of the two-parameter filtration), and it is also small, so it is expected to die quickly w.r.t. the Čech radius parameter. The second cycle is bigger, and slightly less dense, so we can expect it to appear later and survive more on the right side, and the same goes for the third one. The fourth one appears when the Čech radius is large enough (in order to connect the three previous cycles together), and the condensity parameter is large enough as well (such that the three previous cycles are visible).
Figure 19: (Left) The point cloud dataset, colored with the (estimated) density of the sampling. (Right) The candidate decomposition produced by MMA. One can see that four interval summands clearly stand out, and can be interpreted as described in the text. The summands that are induced by noise are all located on the rainbow strip on the left side.
@@
In our next two experiments, we focus on two real-world data sets of point clouds. The first ones, called LargeHypoxicRegion, were obtained from immunohistochemistry in (58). These are made of a few thousand points, each representing a single cell. The others were obtained from applying time-delay embedding in $\mathbb{R}^{2}$ on time series taken from a few data sets (Wine, Plane, OliveOil, Coffee) from the UCR archive (25). On all of these data sets, we computed bi-filtrations using the standard Vietoris-Rips filtration, and the superlevel sets of a Gaussian kernel density estimation (with bandwidth parameter $0.1d$ where $d$ is the diameter of the dataset), and we applied MMA and its competitors on the corresponding multi-parameter persistence modules in homology dimensions 0 and $1$ (note that the ERS can only be computed in degree 0). A typical example of an MMA representation is given in Figure 20.
In our third experiment, we measure the dependence on the number of filtrations using a synthetic data set obtained by sampling $300$ points in the unit square $0,1]^{2}$, computing their Alpha simplicial complex, and generating $n$ random filtration values on the points.
Finally, it is worth noting that we used multi-parameter edge collapses (3) in order to simplify the multi-parameter persistence modules (without losing information) as much as possible before applying Rivet and MMA. The timing of this simplifications are not taken into account, but they are not the computational bottleneck of our computations. Furthermore, for the LargeHypoxicRegion, we thresholded the Rips edges at $0.02$, leading to simplicial complexes of $\sim\!85k$ and $\sim\!125k$ simplices respectively, after simplifications.
Figure 20: (left) LargeHypoxicRegion2, colored with a kernel density estimation of bandwidth $0.01$, (middle, resp. right) Interval decomposition of degree 0 (resp. 1) homology given by MMA. There are $\approx 28k$ non-trivial intervals, each one having a unique color.
@@
We first empirically validate Propositions 3, 4 and 5 by measuring how far is the output of MMA from the data when the precision parameter $\delta$ decreases and the number of lines in $L$ increases. For this, we used the bottleneck distances between the fibered barcodes (on $100$ random diagonal lines) of the outputs of MMA and the ones of the underlying modules as a proxy for the interleaving distances (since they are practically very difficult to evaluate). We call this the estimated matching distance. Results are displayed in Figure 21. One can see that the convergence is empirically linear with the number of lines $|L|$, which is in line with Propositions 3, 4 and 5 (since $|L|$ increases linearly as $\delta$ decreases for a fixed $n$). Note that the distance even reaches 0 on a few cases.
@@
We now compare the running times of MMA with those of Rivet and the ERS. Results can be found in Table 1. It is worth noting that on several occasions, Rivet and the ERS could not produce outputs in reasonable time, due (among other things) to large memory consumption. On the other hands, the fact that MMA produces discretely presented intervals allows to encode them in a sparse manner with their corners. Note that computations with Rivet can be also approximated by coarsening the filtration values, and thus the module. In practice, this corresponds to restricting the 2-module to a $\kappa=\kappa_{x}\times\kappa_{y}$ grid, where $\kappa_{x},\kappa_{y}\in\mathbb{N}$ are the resolutions in both axes, see (47, Section 8.2). The parameters $\kappa$ and $|L|$ have (roughly) the same role, and can be related with $\sqrt{\kappa}\simeq|L|$ (for a given, prescribed precision). Overall, we find that the running times of MMA outperforms its competitors, except when $\kappa$ is very small. However, in this case, the corresponding output of Rivet is very crude as the module is restricted on just a few points, whereas MMA produces intervals that are accurate along whole straight lines: we find that for large data sets, MMA is the only method that can produce accurate representations.
@@
Coffee $0.21\pm 0.01,0.18\pm 0.01$ $9.75\pm 5.92,0.35\pm 0.12$ $0.95\pm 0.56$ $34\pm 0.04 0.0$$093\pm 0.001 0.0$$024\pm 0.001$ $0.16\pm 0.008$
Plane $0.19\pm 0.005,0.18\pm 0.03$ $4.36\pm 2.24,0.28\pm 0.04$ $33.3\pm 17.5,0.56\pm 0.17$ $09\pm 0.03 0.0$$04\pm 0.0 0.0$$12\pm 0.001 0.0$$095\pm 0.004$
Wine $0.21\pm 0.003,0.19\pm 0.007$ $8.50\pm 2.00,0.22\pm 0.01$ $8\pm 0.023 0.22±0$$2\pm 0.04$ $004\pm 0.0 0.0$$016\pm 0.0$ $0.129\pm 0.002$
OliveOil $0.21\pm 0.004,0.19\pm 0.002$ $5.55\pm 1.20,0.31\pm 0.016$ $0.82\pm 0.17$ $1.39\pm 0.03$ $026\pm 0.001 0.0$$058\pm 0.001$ $0.37\pm 0.006$
Worms $0.29\pm 0.082,0.23\pm 0.23$ $19.9\pm 14.4,4.60\pm 5.0$ $31.36\pm 36.24$ $3.85\pm 0.1$ $0.22\pm 0.11$ $0.34\pm 0.15$ $1.35\pm 0.41$
LargeHypoxicRegion1 $1.73,2.88$ ,234--,234 ,--,-- -- $26.4$ $26.6$ $59.4$
LargeHypoxicRegion2 $2.39,6.04$ ,--,-- ,--,-- -- $57.3$ $54.3$ $102.9$
@@
Interestingly, computing 0-dimensional homology is sometimes slower than 1-dimensional homology for Rivet; as it relies on computing minimal free presentations, we think that this comes from the fact that minimal presentations in homology dimension 0 can be more complex than their counterparts in homology dimension 1, i.e., they have much more generators. We also investigate how running times of MMA depend on the number of lines. Unsurprisingly, one can see from Figure 22 that running time increases with the number of lines. However, the dependency looks empirically sublinear, which could come from the fact that even though there are more lines, these lines are closer to each other, and thus matching them with vineyards requires fewer computation steps. This is also highlighted by the running times of LargeHypoxicRegion2, Table 1 which are smaller when computing it over 10001~000 lines than $100$ lines.
Figure 22: Running time (s) needed to run MMA on the UCR datasets.
Figure 23: Running time (s) of MMA w.r.t. the number $n$ of filtrations.
@@
Finally, we investigate how the running times of MMA depend on the number $n$ of filtrations. Although most of the approaches in the literature are limited to $n=2$ parameters, one can see from Figure 23 that MMA can produce outputs in a reasonable amount of time for up to $n\simeq 10$ parameter filtrations. As expected from the complexity of MMA in Equation (9), the running times increase exponentially with $n$.
@@ 8 Conclusion
In this article, we present MMA: a new algorithm for computing topological descriptors for multi-parameter persistence modules taking the form of candidate decompositions. Our algorithm has complexity and running time that are controlled by user-defined parameters, and enjoy several approximation and stability properties. We also showcased the performances of MMA on synthetic and real data sets. Our code is part of the multipers library (50) and is publicly available at https://github.com/DavidLapous/multipers.
Along the way, we identified several directions for future work.
@@
    While the outputs of MMA satisfy approximation guarantees when computed with compatible matching functions, they remain arbitrary to some extent, as discussed at the end of related work and Figure 8.
    We hypothesize that, in the general case, the family $\mathcal{F}$ of approximate decompositions obtained from MMA by varying $\sigma$ across an appropriate family of compatible matching functions $\Sigma$, i.e., $\mathcal{F}=\{\texttt{MMA}{}({\mathbb{M}},L,\sigma)\}_{\sigma\in\Sigma,\delta>0}$, is a complete topological invariant of the module $\mathbb{M}$.
@@
    Practically speaking, the existence and construction of compatible matching functions for $n$-parameter persistence modules with $n>2$ is an open question. We hypothesize that matching functions computed from tracking representative cycles, in a similar way than the construction of the graphcode (39, 40), could provide a step towards that direction. Another possibility includes designing a convex optimization problem that would converge to compatible (and potentially stable) matching functions.
Statements and Declarations
Ethical Approval
Not applicable.
Competing interests
No competing interests.
Authors’ contributions
D.L. M.C. and A.B worked out the proofs and wrote the manuscript, D.L. did the numerical experiments. All authors reviewed the manuscript.
Funding
D.L. was funded by the French government through the 3IA Côte d’Azur Investments, ANR-19-P3IA-0002. M.C. was supported by Agence Nationale de la Recherche through ANR JCJC TopModel ANR-23-CE23-0014, and by the French government, through the 3IA Cote d’Azur Investments in the project managed by the National Research Agency (ANR) with the reference number ANR-23-IACL-0001.
Availability of data and materials.
All of the non-synthetic data sets used are publicly available, and listed below.
  1. 1.
    The immunohistochemistry datasets can be retrieved from
    https://github.com/MultiparameterTDAHistology/SpatialPatterningOfImmuneCells, and
  2. 2.
    The time series datasets can be retrieved from http://www.timeseriesclassification.com/dataset.php.
All of the code used is publicly available as well.
  1. 1.
    Our code, in multipers, at https://github.com/DavidLapous/multipers,
  2. 2.